# Transaction Fee Market Design for Parallel Execution

### v1

@@ Appendix B Proofs Omitted From Section 4
Proof.
Consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\}$ where $\textit{tx}_{1}\simeq(1,\{r_{1}\})$, $\textit{tx}_{2}\simeq(1,\{r_{1}\})$, and $\textit{tx}_{3}\simeq(1,\{r_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\sum_{\textit{tx}\in T}\textsf{gas}_{T}^{B}(\textit{tx})=\frac{3}{4}+\frac{3}{4}+\frac{1}{4}=\frac{7}{4}\neq 2=v(T)$, violating Efficiency. ∎
Appendix C Proofs Omitted From Section 5
C.1. Transaction Bundling (Property 5)
See 5.3
Proof.
Consider the set of transactions $T=\{\textit{tx}_{4}\}$ and three other transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, where $\textit{tx}_{1}\simeq(1,\{r_{2}\})$, $\textit{tx}_{2}\simeq(1,\{r_{2}\})$, $\textit{tx}_{3}\simeq(2,\{r_{2}\})$, and $\textit{tx}_{4}\simeq(1,\{r_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@ C.1. Transaction Bundling (Property 5)
Thus, $\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{2})=\frac{10}{6}>\frac{3}{2}=\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{S}(\textit{tx}_{3})$, which violates Transaction Bundling. ∎
See 5.4
Proof.
@@
$\leq\frac{1}{2^{|T\cup\{\textit{tx}_{3}\}|-1}}\sum_{S\subseteq T}\left[v(S\cup\{\textit{tx}_{3}\})-v(S)\right]\quad=\quad\textsf{gas}^{B}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).\qed$
See 5.5
Proof.
@@
$\leq\frac{t(\textit{tx}_{3})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{3}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{3}\})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).\qed$
See 5.6
Proof.
Consider the set of transactions $T=\{\textit{tx}_{4}\}$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, where $\textit{tx}_{1}\simeq(1,\{r_{1}\})$, $\textit{tx}_{2}\simeq(1,\{r_{1}\})$, $\textit{tx}_{3}\simeq(2,\{r_{1}\})$, and $\textit{tx}_{4}\simeq(1,\{r_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{2})=2>\frac{3}{2}=\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{\textit{ESM}}(\textit{tx}_{3})$, which violates Transaction Bundling. ∎
See 5.7
Proof.
@@
$\leq\frac{v(T\cup\{\textit{tx}_{3}\})}{3^{|T\cup\{\textit{tx}_{3}\}|}}=\textsf{gas}^{\textit{XSM}}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).\qed$
C.2. Set Inclusion (Property 4)
See 5.8
Proof.
Consider the transaction sets $T=\{\textit{tx}_{1}\}$, $T_{1}=\{\textit{tx}_{2},\textit{tx}_{3}\}$, and $T_{2}=\{\textit{tx}_{2},\textit{tx}_{3},\textit{tx}_{4}\}$ where $\textit{tx}_{1}\simeq(1,\{r_{1}\})$, $\textit{tx}_{2}\simeq(1,\{r_{2}\})$, $\textit{tx}_{3}\simeq(1,\{r_{2}\})$, and $\textit{tx}_{4}\simeq(1,\{r_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@ C.2. Set Inclusion (Property 4)
Thus, $\textsf{gas}_{T\cup T_{1}}^{S}(T_{1})=\frac{10}{6}>\frac{36}{24}=\textsf{gas}_{T\cup T_{2}}^{S}(T_{2})$, which violates Set Inclusion. ∎
See 5.9
Proof.
Consider the transaction sets $T=\emptyset$, $T_{1}=\{\textit{tx}_{1},\textit{tx}_{2}\}$, and $T_{2}=\{\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\}$ where $\textit{tx}_{1}\simeq(1,\{r_{1}\})$, $\textit{tx}_{2}\simeq(1,\{r_{1}\})$, and $\textit{tx}_{3}\simeq(1,\{r_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\textsf{gas}_{T\cup T_{1}}^{B}(T_{1})=2>\frac{7}{4}=\textsf{gas}_{T\cup T_{2}}^{B}(T_{2})$, which violates Set Inclusion. ∎
See 5.10
Proof.
@@
$\leq\frac{\sum_{\textit{tx}\in T_{2}}t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T\cup T_{2}}t(\textit{tx}^{\prime})}\cdot v(T\cup T_{2})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup T_{2}}(T_{2}).\qed$
See 5.11
Proof.
@@
$\leq\frac{|T_{2}|}{|T\cup T_{2}|}\cdot v(T\cup T_{2})\quad=\quad\textsf{gas}^{\textit{ESM}}_{T\cup T_{2}}(T_{2}).\qed$
See 5.12
Proof.
Consider the transaction sets $T=\emptyset$, $T_{1}=\{\textit{tx}_{1}\}$, and $T_{2}=\{\textit{tx}_{1},\textit{tx}_{2}\}$, where $\textit{tx}_{1}\simeq(1,\{r_{1}\})$ and $\textit{tx}_{2}\simeq(1,\{r_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\textsf{gas}_{T\cup T_{1}}^{\textit{XSM}}(T_{1})=\frac{1}{3}>\frac{2}{9}=\textsf{gas}_{T\cup T_{2}}^{\textit{XSM}}(T_{2})$, which violates Set Inclusion. ∎
C.3. Resource-Time Sensitivity (Property 3) and Strict Time Sensitivity (Property 2)
See 5.13
Proof.
Consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $R_{1}\subseteq R_{2}$. Then, for any scheduler satisfying property (S3):
@@ C.3. Resource-Time Sensitivity (Property 3) and Strict Time Sensitivity (Property 2)
This establishes Resource-Time Sensitivity.
To also get strict Time Sensitivity, assume $t_{1}<t_{2}$ and $R_{1}=R_{2}$ in the above proof. Note that, when $\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})>0$, the function $\frac{t}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})+t}$ is strictly increasing in $t$, guaranteeing that the first inequality above is strict, so $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}^{\textit{TPM}}(\textit{tx}_{1})<\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}^{\textit{TPM}}(\textit{tx}_{2})$, as required. In the degenerate case where $\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})=0$, the fraction becomes 1 in both cases, but this can only happen when $T=\varnothing$, in which case the value $v$ increases strictly because $v(\{\textit{tx}_{1}\})=t_{1}<t_{2}=v(\{\textit{tx}_{2}\})$. Overall, the first inequality above again holds strictly, ensuring that $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}^{\textit{TPM}}(\textit{tx}_{1})<\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}^{\textit{TPM}}(\textit{tx}_{2})$ still holds, as required. Thus, the strict Time Sensitivity property is satisfied. ∎

### v2

@@ Appendix B Proofs Omitted From Section˜4
Proof B.1.
Consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{1}\})$, and $\textit{tx}_{3}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\sum_{\textit{tx}\in T}\textsf{gas}_{T}^{B}(\textit{tx})=\frac{3}{4}+\frac{3}{4}+\frac{1}{4}=\frac{7}{4}\neq 2=v(T)$, violating Efficiency.
Appendix C Proofs Omitted From Section˜5
C.1 Transaction Bundling (Property˜5)
See 5.5
Proof C.1.
Consider the set of transactions $T=\{\textit{tx}_{4}\}$ and three other transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, where $\textit{tx}_{1}\simeq(1,\{k_{2}\})$, $\textit{tx}_{2}\simeq(1,\{k_{2}\})$, $\textit{tx}_{3}\simeq(2,\{k_{2}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@ C.1 Transaction Bundling (Property˜5)
Thus, $\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{2})=\frac{10}{6}>\frac{3}{2}=\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{S}(\textit{tx}_{3})$, which violates Transaction Bundling.
See 5.6
Proof C.2.
@@
$\leq\frac{1}{2^{|T\cup\{\textit{tx}_{3}\}|-1}}\sum_{S\subseteq T}\left[v(S\cup\{\textit{tx}_{3}\})-v(S)\right]\quad=\quad\textsf{gas}^{B}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
See 5.7
Proof C.3.
@@
$\leq\frac{t(\textit{tx}_{3})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{3}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{3}\})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
See 5.8
Proof C.4.
Consider the set of transactions $T=\{\textit{tx}_{4}\}$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{1}\})$, $\textit{tx}_{3}\simeq(2,\{k_{1}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{2})=2>\frac{3}{2}=\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{\textit{ESM}}(\textit{tx}_{3})$, which violates Transaction Bundling.
See 5.9
Proof C.5.
@@
$\leq\frac{v(T\cup\{\textit{tx}_{3}\})}{3^{|T\cup\{\textit{tx}_{3}\}|}}=\textsf{gas}^{\textit{XSM}}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
C.2 Set Inclusion (Property˜4)
See 5.10
Proof C.6.
Consider the transaction sets $T=\{\textit{tx}_{1}\}$, $T_{1}=\{\textit{tx}_{2},\textit{tx}_{3}\}$, and $T_{2}=\{\textit{tx}_{2},\textit{tx}_{3},\textit{tx}_{4}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{2}\})$, $\textit{tx}_{3}\simeq(1,\{k_{2}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@ C.2 Set Inclusion (Property˜4)
Thus, $\textsf{gas}_{T\cup T_{1}}^{S}(T_{1})=\frac{10}{6}>\frac{36}{24}=\textsf{gas}_{T\cup T_{2}}^{S}(T_{2})$, which violates Set Inclusion.
See 5.11
Proof C.7.
Consider the transaction sets $T=\emptyset$, $T_{1}=\{\textit{tx}_{1},\textit{tx}_{2}\}$, and $T_{2}=\{\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{1}\})$, and $\textit{tx}_{3}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\textsf{gas}_{T\cup T_{1}}^{B}(T_{1})=2>\frac{7}{4}=\textsf{gas}_{T\cup T_{2}}^{B}(T_{2})$, which violates Set Inclusion.
See 5.12
Proof C.8.
@@
$\leq\frac{\sum_{\textit{tx}\in T_{2}}t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T\cup T_{2}}t(\textit{tx}^{\prime})}\cdot v(T\cup T_{2})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup T_{2}}(T_{2}).$
See 5.13
Proof C.9.
@@
$\leq\frac{|T_{2}|}{|T\cup T_{2}|}\cdot v(T\cup T_{2})\quad=\quad\textsf{gas}^{\textit{ESM}}_{T\cup T_{2}}(T_{2}).$
See 5.14
Proof C.10.
Consider the transaction sets $T=\emptyset$, $T_{1}=\{\textit{tx}_{1}\}$, and $T_{2}=\{\textit{tx}_{1},\textit{tx}_{2}\}$, where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$ and $\textit{tx}_{2}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
@@
Thus, $\textsf{gas}_{T\cup T_{1}}^{\textit{XSM}}(T_{1})=\frac{1}{3}>\frac{2}{9}=\textsf{gas}_{T\cup T_{2}}^{\textit{XSM}}(T_{2})$, which violates Set Inclusion.
C.3 Storage Key-Time Monotonicity (Property˜3) and Strict Time Monotonicity (Property˜2)
See 5.15
Proof C.11.
Consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$. Then, for any scheduler satisfying property (S3):
@@ C.3 Storage Key-Time Monotonicity (Property˜3) and Strict Time Monotonicity (Property˜2)
This establishes Storage Key-Time Monotonicity.
To also get strict Time Monotonicity, assume $t_{1}<t_{2}$ and $K_{1}=K_{2}$ in the above proof. Note that, when $\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})>0$, the function $\frac{t}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})+t}$ is strictly increasing in $t$, guaranteeing that the first inequality above is strict, so $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}^{\textit{TPM}}(\textit{tx}_{1})<\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}^{\textit{TPM}}(\textit{tx}_{2})$, as required. In the degenerate case where $\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})=0$, the fraction becomes 1 in both cases, but this can only happen when $T=\varnothing$, in which case the value $v$ increases strictly because $v(\{\textit{tx}_{1}\})=t_{1}<t_{2}=v(\{\textit{tx}_{2}\})$. Overall, the first inequality above again holds strictly, ensuring that $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}^{\textit{TPM}}(\textit{tx}_{1})<\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}^{\textit{TPM}}(\textit{tx}_{2})$ still holds, as required. Thus, the strict Time Monotonicity property is satisfied.